You need 880 mL of a 75% alcohol solution. On hand, you have a 20% alcohol mixture. How much of the 20% alcohol mixture and pure alcohol will you need to obtain the desired solution?
You will need
mL of the 20% solution
and
mL of pure alcohol.

Answer by stanbon(75887)

You need 880 mL of a 75% alcohol solution. On hand, you have a 20% alcohol mixture. How much of the 20% alcohol mixture and pure alcohol will you need to obtain the desired solution?
--------------------------------------
alcohol + alcohol = alcohol
0.20x + 1(880-x) = 0.75*880
Multiply thru by 100 to get:
20x + 100*880 -100x = 75*880
-80x = - -20*880
x = 220 mml ( amt. of pure alcohol needed for the mixture)
880-220 = 660mml ( amt. of 20% solution you need for the mixture)
===========
Cheers,
Stan H.